Find the value of a such that the line y=x^2+a has exactly one intersection point with the parabola with equation y=x^2+3x+2.
Answer by Alan3354(69443)

Find the value of a such that the line y=x^2+a has exactly one intersection point with the parabola with equation y=x^2+3x+2.
--------------------
I see only one intersection of the 2 parabolas for any value of a.
You said "line y=x^2+a", but that's a parabola. Did you mean y = x+a?